A mass of 5kg is on top of a building of height 10m. Calculate its potential energy.
The mass then falls at a speed of 8 m/s in 4 sec, find its KE and momentum p. What is the Force and impulse the mass transfers to the ground?
1
Expert's answer
2019-11-24T16:52:51-0500

Potential energy:


"E=mgh=(5)(9.8)(10)=490\ J"


"KE=0.5mv^2=0.5(5)(8)^2=160\ J"

"P=mv=(5)(8)=40\frac{kgm}{s}"

​ ​ The impulse the mass transfers to the ground:


"I=P=40\frac{kgm}{s}"

The force:


"F=\frac{I}{t}=\frac{40}{4}=10\ N"